OpenStudy (anonymous):

What are the steps to solve this problem? 0.000027/0.000009 =

OpenStudy (ranga):

Multiply both numerator and denominator by: \[\Large 10^{6}\]

OpenStudy (anonymous):

It's division, not a fraction

OpenStudy (ranga):

\[\frac{ 0.000027 }{0.000009 } = \frac{ 0.000027 }{0.000009 } \times \frac{ 10^{6} }{10^{6}} = \frac{ 27 }{ 9 } = 3\]